Mr. Bacani borrowed money from the bank. He received from the bank P1,842 and promised to repay P2,000 at the end of 10 months. Determine the rate of simple interest.


1
Expert's answer
2022-03-08T12:55:14-0500

Let the amount of loan borrowed be (P) = 1842

Let the amount of money to be paid (X) =2000

The payment period (n) = 10 months

The amount of Simple interest (S.I)

(S.I) = 2000-1842 = 158


But simple interest is given by

"(S.I)=(Q*R\/100*T)"


P158="(P1842*R\/100*10)"

P158*100=P18420 "R"

"\therefore" R=0.8577633

The monthly rate of simple interest rate is approximately 0.8578%